In the study of calculus, the difference quotient f(a+h)-f(a)/h is often found and simplified. Find and simplify this quotient for the function f(x). Complete parts a through d.
f(x)=8/x
A. Find f(a+h).
f(a+h)=____________
B. Find f(a)
f(a)= _____________
C. Use steps a and b to find f(a+h)-f(a)/h
f(a+h)-f(a)/h = _______________ (do not simplify)
D. Simplify the result of step c.
f(a+h)-f(a)/h=_______________

Answer by stanbon(75887)

In the study of calculus, the difference quotient f(a+h)-f(a)/h is often found and simplified. Find and simplify this quotient for the function f(x). Complete parts a through d.
f(x)=8/x
A. Find f(a+h).
f(a+h)= 8/(a+h)
----------------------
B. Find f(a)
f(a)= 8/a
----------------------
C. Use steps A and B to find f(a+h)-f(a)/h
f(a+h)-f(a)/h = 8/(a+h)-8/a (do not simplify)
--------------------------
D. Simplify the result of step C.
f(a+h)-f(a)/h= [8a-8(a+h)]/[a(a+h)]
= [-8h]/[a^2+ah]
------
Cheers,
Stan H.
----------